OpenStudy (anonymous):

A student performed a chemical reaction which involved 3 reactants: A, B and C and measured the initial reaction rate but varied initial concentrations. The results showed that doubling the concentration of A doubled the reaction rate, doubling the concentration of B quadrupled the reaction rate and doubling the concentration of C had no effect on the rate of reaction. Find the overall reaction order and write the rate equation.

OpenStudy (anonymous):

Well your rate would be \[R=[A]^x[B]^y[C]^z\] if you double A and R doubles as a result, keeping [B] and [C] constant \[R=[A]^x[B][C] \]\[2R=[2A]^x[B][C]\] and you can equate this with the original by multiplying the entire equation by 2. \[2R=2[A]^x[B][C]\]then we equate the two, since 2R = 2R. \[2[A]^x[B][C] = [2A]^x[B][C]\]the concentrations of B and C will cancel (really just doing a silly long version, typically you'd ignore the other concentrations because they will cancel as in the next step)\[\frac{ 2[A^x][B][C] }{ [B][C] }=\frac{ [2A]^x[B][C] }{ [B][C] }\]\[2[A^x]=[2A^x]\] so now you want a value for the exponent x that makes this equation true, and that'd be 1. So x=1. Do the same for B and C and the sum of your exponents (x + y + z) is the overall order of this rxn.